# Cascade of topological phase transitions and revival of topological zero modes in imperfect double helical liquids

### v1

@@ Abstract
Two parallel helical edge channels hosting interacting electrons, when proximitized by local and nonlocal pairings, can host time-reversal-invariant pairs of topological zero modes at the system corners. Here we show that realistic imperfections substantially enrich the physics of such proximitized double helical liquids. Specifically, we analyze this platform and its fractional counterparts in the presence of pairing and interaction asymmetries between the two channels, as well as random spin-flip terms arising from either magnetic disorder or coexisting charge disorder and external magnetic fields. Using renormalization-group analysis, we determine how Coulomb interactions, pairings, and magnetic disorder collectively influence the transport behavior and topological properties of the double helical liquid. As the system transitions from class DIII to class BDI, an additional topological phase supporting a single Majorana zero mode per corner emerges. We further show how additional pairing or Coulomb asymmetry influences the stability of various topological phases and uncovers a revival of Majorana zero modes and cascades of transitions through topological phases characterized by a $\mathbb{Z}$ invariant, which are accessible through controlling the electrical screening effect. In contrast to conventional understanding, disorder is not merely detrimental, as it in general allows for a tuning knob that qualitatively reshapes the topological superconductivity in imperfect helical liquids.
@@ I Introduction
Among various alternatives, such as atomic chains and other low-dimensional platforms [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30], a promising route for avoiding external magnetic fields is to engineer superconductivity in helical liquids [31, 32, 33, 34]. Helical liquids arise at the edges or hinges of time-reversal-invariant topological insulators [35, 36, 37, 38, 39, 40, 41, 42, 33, 43, 44, 45], providing a magnetic-field-free platform with symmetry-protected one-dimensional modes that are well suited for realizing topological and correlated quantum states. When brought into proximity with superconductors, such helical channels can host Majorana and parafermion zero modes [46, 47, 48, 49, 50, 31, 51, 52, 32, 53, 54, 55, 33, 56, 57, 58, 59], providing time-reversal-invariant settings for stabilizing and detecting zero modes, as well as forming topological quantum bits [60, 61, 62]. Recent developments have further extended the platform of helical liquids to moiré and twisted bilayer systems [63, 64], where fractional analogues of quantum Hall edges can emerge [63]. These systems open new opportunities for realizing correlated and topological phases.
@@
In this work, we investigate a system of two interacting helical channels with both intrachannel and interchannel pairings, and introduce imperfections commonly present in realistic devices, including random spin-flip backscattering as well as pairing and Coulomb asymmetries between the two channels. These ingredients arise naturally from magnetic impurities, in-plane fields combined with charge disorder, and inhomogeneous local environment of the two channels. We derive an analytical expression for the number of topological zero modes as a function of system parameters, and combine it with the renormalization-group (RG) framework. We show that spin-flip backscattering generically detunes the two zero-mode conditions that are degenerate in the clean limit, opening a new phase where the system supports a single Majorana zero mode. We also demonstrate that interactions and disorder can drive realistic systems into this regime even when the clean limit remains topologically trivial. We further find that backscattering enhances the effects of weak pairing asymmetry, leading to a controllable revival of topological zero modes and a cascade of transitions between phases labeled by a $\mathbb{Z}$ invariant. Taken together, these results show how realistic imperfections reorganize both the topological and transport properties of (fractional) helical liquids, with electrically tunable features in the platform.
@@
The rest of this article is organized as follows. In Sec. II, we introduce our setup and the corresponding bosonized Hamiltonian, consisting of interacting electrons in two helical channels with proximity-induced local and nonlocal pairings and random spin-flip backscatterings. In Sec. III, we discuss the RG flow equations and representative RG flow diagrams. In Sec. IV, we explore the transport properties of the system. We use the renormalized couplings to analyze how various phases evolve in Sec. IV.1, and compute the localization length and temperature in the insulating phase in Sec. IV.2. In Sec. V, we characterize the system according to its topological properties and the associate zero modes. We discuss the formula of the number of the zero modes and examine how various ingredients existing in imperfect helical liquids influence the topology in Sec. V.1. Combining with RG analysis, we obtain topological phase diagrams in Sec. V.2. Finally, in Sec. VI, we summarize our main findings and discuss possible extensions and experimental realizations. Technical details are given in the appendices. The experimentally relevant platforms and their material parameters are summarized in Appendix A. The derivation of random spin-flip backscattering terms are given in Appendix B. The derivation of the RG flow equations, including the generalization to fractional helical liquids, is in Appendix C. The symmetry analysis of the effective model, solutions of the corresponding Bogoliubov-de Gennes equation and criterion of the zero modes are collected in Appendix D. Additional details of the numerical analysis are provided in Appendix E.
@@ II Setup and model
where $\phi_{n}$ and $\theta_{n}$ denote the bosonic dual fields with the velocities $u_{n}\equiv v_{{\rm F}}/(K_{n}m)$ and interaction parameter $K_{n}$. In our convention, the value $K_{n}=1$ corresponds to the noninteracting limit, whereas $K_{n}<1$ ($K_{n}>1$) indicates repulsive (attractive) electron-electron interactions.
@@ III RG flow equations and flow diagrams
where we include the leading-order contributions in each equation.
@@ V.1 Characterizing topological phases
It is instructive to compare our findings with previous studies on time-reversal invariant topological superconductors in bilayer systems, such as Ref. [129]. In those setups, the driving mechanism for Majorana corner states is to have a dominant coherent single-particle interlayer tunneling over the pairing [129]. In the absence of interlayer tunneling, however, the clean limit of a $\pi$-junction ($V_{\pm}=0,\,\Delta_{c}=0$) would result in the impossible condition $|\Delta_{-}(l^{*})|<0$ in Eq. (V.1). Thus, a pure $\pi$-phase difference is insufficient to stabilize Majorana zero modes.
@@ V.2.3 Imperfection-induced topological phases
VI Discussion

### v2

@@ Abstract
Two parallel helical edge channels hosting interacting electrons, when proximitized by local and nonlocal pairings, can host time-reversal-invariant pairs of topological zero modes at the system corners. Here we show that realistic imperfections substantially enrich the physics of such proximitized double helical liquids. Specifically, we analyze this platform and its fractional counterparts in the presence of pairing and interaction asymmetries between the two channels, as well as random spin-flip terms arising from either magnetic disorder or coexisting charge disorder and external magnetic fields. Using renormalization-group analysis, we determine how Coulomb interactions, pairings, and magnetic disorder collectively influence the transport behavior and topological properties of the double helical liquid. As the system transitions from class DIII to class BDI, an additional topological phase supporting a single Majorana zero mode per corner emerges. We further show how additional pairing or Coulomb asymmetry influences the stability of various topological phases and uncovers a revival of Majorana zero modes and cascades of transitions through topological phases characterized by a $\mathbb{Z}$ invariant, which are accessible through controlling the electrical screening effect. We also analyze the spatial structure of the zero modes and the bulk gap closing through channel-resolved density profiles. In contrast to conventional understanding, disorder is not merely detrimental, as it in general allows for a tuning knob that qualitatively reshapes the topological superconductivity in imperfect helical liquids.
@@ I Introduction
Among various alternatives, such as atomic chains and other low-dimensional platforms [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31], a promising route for avoiding external magnetic fields is to engineer superconductivity in helical liquids [32, 33, 34, 35, 36]. Helical liquids arise at the edges or hinges of time-reversal-invariant topological insulators [37, 38, 39, 40, 41, 42, 43, 44, 35, 45, 46, 47, 48, 49, 50], providing a platform with time-reversal-invariant one-dimensional modes that are well suited for realizing topological and correlated quantum states. When brought into proximity with superconductors, such helical channels can host Majorana and parafermion zero modes [51, 52, 53, 54, 55, 32, 56, 57, 33, 58, 59, 60, 35, 61, 62, 63, 64], providing time-reversal-invariant settings for stabilizing and detecting zero modes, as well as forming topological quantum bits [65, 66, 67]. Recent developments have further extended the platform of helical liquids to moiré and twisted bilayer systems [68, 69], where fractional analogues of quantum Hall edges can emerge [68]. These systems open new opportunities for realizing correlated and topological phases.
@@
In this work, we investigate a system of two interacting helical channels with both intrachannel and interchannel pairings, and introduce imperfections commonly present in realistic devices, including random spin-flip backscattering as well as pairing and Coulomb asymmetries between the two channels. These ingredients arise naturally from magnetic impurities, in-plane fields combined with charge disorder, and inhomogeneous local environment of the two channels. We derive an analytical expression for the number of topological zero modes as a function of system parameters, and combine it with the renormalization-group (RG) framework. We show that spin-flip backscattering generically detunes the two zero-mode conditions that are degenerate in the clean limit, opening a new phase where the system supports a single Majorana zero mode. We also demonstrate that interactions and disorder can drive realistic systems into this regime even when the clean limit remains topologically trivial. We further find that backscattering enhances the effects of weak pairing asymmetry, leading to a controllable revival of topological zero modes and a cascade of transitions between phases labeled by a $\mathbb{Z}$ invariant. Finally, we analyze the channel-resolved density profiles, which reveal observable features in scanning probes upon varying the system parameters. Taken together, these results show how realistic imperfections reorganize both the topological and transport properties of (fractional) helical liquids, with electrically tunable features in the platform.
@@
The rest of this article is organized as follows. In Sec. II, we introduce our setup and the corresponding bosonized Hamiltonian, consisting of interacting electrons in two helical channels with proximity-induced local and nonlocal pairings and random spin-flip backscatterings. In Sec. III, we discuss the RG flow equations and representative RG flow diagrams. In Sec. IV, we explore the transport properties of the system. We use the renormalized couplings to analyze how various phases evolve in Sec. IV.1, and compute the localization length and temperature in the insulating phase in Sec. IV.2. In Sec. V, we characterize the system according to its topological properties and the associate zero modes. We discuss the formula of the number of the zero modes and examine how various ingredients existing in imperfect helical liquids influence the topology in Sec. V.1. Combining with RG analysis, we obtain topological phase diagrams in Sec. V.2. In Sec. VI, we construct both the transport- and topology-based phase diagrams in three-dimensional parameter spaces. In Sec. VII, we visualize the spatial evolution of the Majorana zero modes and the bulk gap closing using channel-resolved density profiles. Finally, in Sec. VIII, we summarize our main findings and discuss possible extensions and experimental realizations. Technical details are given in the appendices. The experimentally relevant platforms and their material parameters are summarized in Appendix A. The derivation of random spin-flip backscattering terms are given in Appendix B. The derivation of the RG flow equations, including the generalization to fractional helical liquids, is in Appendix C. The symmetry analysis of the effective model, solutions of the corresponding Bogoliubov-de Gennes equation and criterion of the zero modes are collected in Appendix D. Additional details of the numerical analysis are provided in Appendix E.
@@ II Setup and model
where $\phi_{n}$ and $\theta_{n}$ denote the bosonic dual fields with the velocities $u_{n}\equiv v_{{\rm F}}/(K_{n}m)$ and interaction parameter $K_{n}$. In our convention, the value $K_{n}=1$ corresponds to the noninteracting limit, whereas $K_{n}<1$ ($K_{n}>1$) indicates repulsive (attractive) electron-electron interactions. It has been shown that the interaction strength in the edge channels can be electrically tuned by gates [47].
@@ III RG flow equations and flow diagrams
where we include the leading-order contributions.
@@ V.1 Characterizing topological phases
It is instructive to compare our findings with previous studies on time-reversal invariant topological superconductors in bilayer systems, such as Ref. [134]. In those setups, the driving mechanism for Majorana corner states is to have a dominant coherent single-particle interlayer tunneling over the pairing [134]. In contrast, the driving mechanism in Eq. (V.1) is the nonlocal pairing $\Delta_{c}$. When we neither have this term nor intelayer tunneling, the clean limit of a $\pi$-junction ($V_{\pm}=0,\,\Delta_{c}=0$) would result in the impossible condition $|\Delta_{-}(l^{*})|<0$ in Eq. (V.1). Thus, a pure $\pi$-phase difference is insufficient to stabilize Majorana zero modes.
@@ V.2.3 Imperfection-induced topological phases
In the following section, we explore a broader parameter space by constructing three-dimensional phase diagrams that capture the interplay among interactions, disorder effects, and proximity-induced pairing.
VI Phase diagrams in three-dimensional parameter space
In this section, we present three-dimensional phase diagrams that elucidate the intertwined roles of interactions, disorder-induced backscattering, and various pairing processes in helical channels. As above, the phase structure is analyzed from both transport and topological perspectives for complementary representations.
In the first set of phase diagrams, we vary the average interaction strength $K_{+}(0)$, the average local pairing strength $\tilde{\Delta}{{}_{+}}(0)$, and the average backscattering strength $\tilde{D}{{}_{+}}(0)$, thereby highlighting the competition between local and nonlocal pairing processes in the presence of interactions and disorder. In the second set, $\tilde{\Delta}{{}_{+}}(0)$ is replaced by the pairing asymmetry $\tilde{\Delta}_{-}(0)$, allowing us to directly isolate and visualize the impact of pairing imbalance on the phase structure. This formulation provides a unified view of how interactions, disorder, and pairing asymmetry cooperate to shape transport behavior and topological stability.
Figure 10: Transport-based phase diagrams in three-dimensional parameter space, including (a–b) local-to-nonlocal pairing ratio $\Delta_{+}(0)/\Delta_{c}(0)$ and (c) pairing asymmetry $\tilde{\Delta}{{}_{-}}(0)$. The regions corresponding to the insulating phase (blue) and the local superconducting phase in channel 2 (orange) or channels 1 and 2 (red) are rendered with increased transparency to reveal the internal structure. The adopted parameter values include (a) $\tilde{\Delta}_{-}(0)=0$, (b) $\tilde{\Delta}_{-}(0)=0.01$, and (c) $\tilde{\Delta}_{+}(0)=0.03$. In each case, we set $\tilde{\Delta}_{c}(0)=0.01$, and $K_{-}(0)=\tilde{D}_{-}(0)=0$. See Table 4 for the complete set of the adopted parameter values.
We first discuss the phase diagrams with a varying average local pairing strength, as shown in Fig. 10(a–b). We observe that increasing the local pairing strength suppresses the nonlocal superconducting phase in both symmetric and asymmetric cases, as displayed in Fig. 10(a) and Fig. 10(b), respectively. Additionally, the asymmetry introduces distinct features in Fig. 10(b). Specifically, we observe the emergence of the asymmetric local superconducting phase at weak interactions and an asymmetric insulating phase (light blue), where one channel localizes faster than the other.
The corresponding topological phase diagrams in Fig. 11(a–b) confirm the robustness of the $N_{\rm mzm}=2$ phase at a weak backscattering when the nonlocal pairing dominates. In the symmetric case in Fig. 11(a), increasing backscattering strength eventually destroys the Majorana zero modes. Interestingly, the three-dimensional phase diagrams reveal that the $N_{\mathrm{mzm}}=1$ phase emerges over a broad region of parameter space. As illustrated in the asymmetric case in Fig. 11(b), a robust $N_{\mathrm{mzm}}=1$ region appears between the $N_{\mathrm{mzm}}=2$ phases and remains accessible in a wide range of backscattering strengths. As the local pairing strength increases, this intermediate phase eventually shrinks. The three-dimensional plots clearly reveal that multiple phase transitions and the revival of zero modes can be induced by tuning a single control parameter, such as the backscattering strength (via external magnetic fields), the interaction strength (through screening effects), or the balance between local and nonlocal pairing (controlled by channel separation or the proximity interface).
Next, we discuss the transport-based phase diagram in Fig. 10(c), which incorporates pairing asymmetry. Consistent with previous observations, we note that increasing pairing asymmetry shifts the system toward regions where asymmetric phases are favored, such as the asymmetric local superconducting phases at weak interactions, and in some regions, asymmetric insulating phases, where one of the channels localizes faster.
The corresponding topological phase diagram with pairing asymmetry is shown in Fig. 11(c), complementing Figs. 11(a–b). Consistent with the analysis in the previous section, the $N_{\mathrm{mzm}}=1$ phase emerges immediately upon deviating from the perfectly symmetric limit $\tilde{\Delta}_{-}(0)=0$ and remains accessible over a broad range of backscattering strengths. This has important experimental implications, as pairing imbalance between channels and weak random spin-flip backscattering are ubiquitous in realistic systems.
Finally, we briefly remark on our findings in view of the $\pi$-junction limit discussed in Sec. V. An ideal $\pi$-junction corresponds to the limit in which the local pairings in the two channels have opposite signs, thus $|\Delta_{-}(0)/\Delta_{+}(0)|\to\infty$. Our analysis shows that the presence of a residual uniform component $\Delta_{+}$ can stabilize multiple phases, revealing rich behavior in nonideal $\pi$ junctions. Moreover, the persistence of the topological phase in Fig. 11(c) in the presence of disorder indicates that $\pi$-junction-based setups can have robust topological phases against realistic backscattering effects.
VII Microscopic features of the Majorana zero modes
In this section, we examine how the Majorana zero modes evolve through the series of topological phase transitions and discuss their features in local probes. To this end, we focus on the representative parameter sets near the point $P$ marked in Fig. 7(b). In this region, tuning either the electron-electron interaction $K_{+}(0)$ or the backscattering $\tilde{D}_{+}(0)$ strength alone results in the phase transitions with the sequence $N_{\rm mzm}=2\rightarrow 1\rightarrow 2$. The scanned ranges of the $K_{+}(0)$ and $\tilde{D}_{+}(0)$ values used in Figs. 12(a–b) were chosen to illustrate this path, where one of the Majorana zero modes is destroyed and then reappears, signifying a tunable revival.
Figure 11: Topological phase diagrams in three-dimensional parameter space, including (a–b) local-to-nonlocal pairing ratio $\Delta_{+}(0)/\Delta_{c}(0)$ and (c) pairing asymmetry $\tilde{\Delta}{{}_{-}}(0)$. The adopted parameter values include (a) $\tilde{\Delta}_{-}(0)=0$, (b) $\tilde{\Delta}_{-}(0)=0.01$, and (c) $\tilde{\Delta}_{+}(0)=0.03$. In each case, $\tilde{\Delta}_{c}(0)=0.01$, and $K_{-}(0)=\tilde{D}_{-}(0)=0$. See Table 4 for the complete set of the adopted parameter values.
To visualize how the modes evolve, we compute their corresponding wavefunctions following the algebra in Appendix D, with the renormalized couplings acquired from the above RG analysis. After obtaining the zero-energy orthonormal modes, $\Phi_{\rm mzm,1}(r)$ and $\Phi_{\rm mzm,2}(r)$, we reconstruct the full wavefunctions by reinstalling the fast-oscillating factors $e^{\pm ik_{F}r}$,
$\Psi_{\text{mzm},j}(r)=e^{ik_{F}r\eta^{z}\sigma^{z}}\Phi_{\text{mzm},j}(r),$ (18)
expressed in the physical basis $(\psi_{1\downarrow},\psi_{1\uparrow},\psi_{2\downarrow},\psi_{2\uparrow},\psi_{1\downarrow}^{\dagger},\psi_{1\uparrow}^{\dagger},\psi_{2\downarrow}^{\dagger},\psi_{2\uparrow}^{\dagger})^{T}$. To proceed, we define
$\rho_{j}(r)\equiv|\Psi_{\text{mzm},j}(r)|^{2}.$ (19)
for the $j$th zero mode, with $j\in\{1,2\}$. Since the Majorana wavefunctions $\Psi_{\text{mzm},j}(r)$ satisfy self-conjugation, the electron and hole components contribute equally. Consequently, $\rho_{j}(r)$ corresponds to twice the particle component of the density profiles. Notably, the wavefunction has finite support in both channels, reflecting the composite nature of the zero modes discussed in Appendix D.
Motivated by the local scanning probes with high spatial resolution, we examine the channel-resolved density profiles. To this end, we decompose $\rho_{j}(r)$ into $\rho_{j}^{(n)}(r)\equiv|\Psi^{(n)}_{\text{mzm},j}(r)|^{2}$ by projecting
$\Psi_{\text{mzm},j}^{(1,2)}(r)\equiv\left[\eta^{0}\left(\frac{\tau^{0}\pm\tau^{z}}{2}\right)\sigma^{0}\right]\Psi_{\text{mzm},j}(r),$ (20)
and evaluate them using the RG results.
To visualize the evolution of the zero modes accross the topological phase transitions, Figs. 12(a,b) display the spatial density maps as functions of the interaction parameters $K_{+}(0)$ and backscattering strengths $\tilde{D}_{+}(0)$. Here, we plot the total density in each channel, $\sum_{j}\rho_{j}^{(n)}(r)$, combining the density of all the zero modes present in the system. In contrast to nonhelical systems [13], the resulting density profiles here do not exhibit spatial oscillations, reflecting the underlying helical nature of the channels.
We first discuss Fig. 12(a), which tracks the density evolution as the interaction parameter $K_{+}(0)$ is varied. Throughout this range, the renormalized nonlocal pairing $\tilde{\Delta}_{c}(l^{*})$ remains the dominant scale, while the local pairings $\tilde{\Delta}_{1,2}(l^{*})$ gradually increase and the backscattering strengths $\tilde{D}_{1,2}(l^{*})$ gradually decrease as $K_{+}(0)$ is raised. At stronger interactions, the topological criterion in Eq. (V.1) is satisfied for both $\varepsilon=+$ and $\varepsilon=$-, and the system hosts two well-localized zero modes around $r=0$. On the $r<0$ side, where the nonlocal pairing is absent, the modes appear more extended. This follows directly from the form of the decay constants on that side: as given in Eq. (69), $\kappa_{n,\varepsilon}^{<}=\big|\Delta_{n}+\varepsilon V_{n}\big|/(\hbar v_{F})$, which depends only on the local channel parameters. By contrast, on the $r>0$ side the decay constants $\kappa_{\lambda,\varepsilon}^{>}=\left|(\Delta_{+}+\varepsilon V_{+})+\lambda\sqrt{(\Delta_{-}+\varepsilon V_{-})^{2}+\Delta_{c}^{2}}\right|/(\hbar v_{F})$ are enhanced by the presence of $\Delta_{c}$, leading to sharper localization. As the interactions weaken (middle region), only one of the two zero-mode conditions remains satisfied, leaving a single Majorana zero mode, in line with the behavior described in Sec. V.2.1. Further weakening the interactions (upper region) restores the second zero-mode condition, leading to the reappearance of the second topological zero mode and the enhanced total density. At the same time, the profiles broaden compared to the lower region, and are less localized. This behavior reflects the changing balance between the local and nonlocal pairings. Namely, as the interactions weaken further, the RG flow drives the backscattering strengths $\tilde{D}_{1,2}(l^{*})$ closer to zero, so the presence of the zero modes is set almost entirely by the pairings. In this regime, the criterion for $N_{\rm mzm}$ approaches the familiar condition $\Delta_{c}^{2}>\Delta_{1}\Delta_{2}$ in Ref. [33]. Eventually the local pairings become comparable to the nonlocal one, and the Majorana zero modes extend over longer distances and eventually vanish.
In Fig. 12(b), we vary the backscattering strength $\tilde{D}_{+}(0)$ to trace a vertical path, with $N_{\rm mzm}=2\to 1\;(\varepsilon=+)\to 2\to 1\;(\varepsilon=-)$ in Fig. 7(b). Deep inside the topological phase centered near $\tilde{D}_{+}(0)\sim 10^{-5}$, the zero modes exhibit their sharpest localization and highest intensity at $r=0$. However, as $\tilde{D}_{+}(0)$ approaches the phase boundaries on either side, the confinement weakens as the gap closes and the $r=0$ intensity fades.
Figure 12: Channel-resolved total density profiles $\sum_{j}\rho_{j}^{(n)}(r)$ for channel $n=1$ (left panels) and $n=2$ (right panels). (a) Density evolution as a function of $K_{+}(0)$ with fixed $\tilde{D}_{+}(0)=10^{-6}$. (b) Density evolution as a function of $\tilde{D}_{+}(0)$ with fixed $K_{+}(0)=0.6$. The remaining fixed parameters correspond to point $P$ in Fig. 7(b), $\tilde{\Delta}_{+}(0)=0.03$, $\tilde{\Delta}_{-}(0)=0.01$, $\tilde{\Delta}_{c}(0)=0.01$, and $\tilde{D}_{-}(0)=K_{-}(0)=0$. See Table 4 for the complete set of the adopted parameter values.
As mentioned above, the density maps provide a qualitative indication of the bulk gap. As derived in Eqs. (69) and (72), the spatial decay constants $\kappa$ are directly proportional to the effective gaps on the respective sides of the interface. Consequently, regions of high intensity near $r=0$ signify strong confinement and a robust topological gap. Conversely, near a phase transition where $N_{\rm mzm}$ changes, the gap closes, driving the decay constant to zero, thus serving as a visual signature of the bulk gap closing.
Accompanying the zero-energy spectroscopic signatures [135], the spatial structure of the Majorana zero modes shown in Fig. 12 can be accessed using scanning tunneling microscopy. Deep inside a topological phase, sharply localized modes are expected near the system corners ($r=0,L$). In contrast, broader spatial profiles are anticipated as the system approaches phase boundaries.
Beyond the spatial density, the internal structure of these wavefunction relates to broader discussions in the existing literature. Ref. [136] retained only the electronic component when defining spin polarization, focusing on quantities accessible via spin-polarized STM, while Ref. [137] incorporated both electron and hole contributions in the spin polarization to account for the electron-hole overlap effects. Ref. [138] focused exclusively on Majorana polarization, which is intrinsically defined using both particle and hole degrees of freedom. Importantly, these works primarily analyze bulk states at finite energies rather than the topological zero modes on which we focus here.
We mention that the density profiles displayed in Fig. 12 correspond to contributions only from the zero modes. Experimental probes, however, may effectively integrate contributions from states away from $E=0$. As a result, a complementary analysis of finite-energy bulk states is necessary to provide additional context, such as background contrast, for experimental characterization. Finally, distinguishing topological Majorana zero modes from trivial bound states remains an important experimental challenge. Other indicators could be useful, including Majorana polarization [136, 138], which quantifies particle-hole overlap, as well as the sign reversal of the spin polarization of bulk states across a topological phase transition [137].
VIII Discussion